# Preference Optimization by Estimating the Ratio of the Data Distribution

### v1

@@ A.1 Proof of Proposition 1
From the Bradley-Terry model we discussed in Eq. 5, the preference data distribution $p_{\text{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}|{\mathbf{x}})$ can be expressed as
@@ A.2 Proof of Theorem 2
Let $\operatorname*{arg\,min}_{\pi_{\theta}}D_{h}(R_{\text{data}}||R_{\theta})=\pi_{\hat{\theta}}$. We have followings from Eqs. 6 and 9:
@@
From the property of Bregman divergence [7], $B_{h}\left(R_{\text{data}}({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})||R_{\theta}({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})\right)$ achieves its minimum value of 0 if and only if $R_{\text{data}}({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})=R_{\theta}({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$ for a given point $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$. Under the assumption that the data distribution is fully supported, the minimizer $\pi_{\hat{\theta}}$ is guaranteed to satisfy $R_{\text{data}}=R_{\hat{\theta}}$ for all $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$. From the definition of $R_{\text{data}}$ and $R_{\theta}$ in Eq. 8, we obtain the following:
@@
From the definition of $\pi^{*}$ in Proposition 1, we obtain the following relation, which holds for all points $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$.
@@ A.5 BPO recovers DPO as a special case
This section shows that BPO includes DPO as a special case as discussed in Section 3.3. Consider $h(R)=\frac{R\log{R}-(1+R)\log{(1+R)}}{2}$ which corresponds to the logistic regression in Table 2, which has a derivative function as $h^{\prime}(R)=\frac{\log{R}-\log{(1+R)}}{2}$ and the expression inside the expectation operator of $\mathcal{L}^{h}_{\textrm{BPO}}(R_{\theta};p_{\textrm{data}})$ in Eq. 10 becomes:
@@ A.6 BPO with $R_{\theta}^{f\text{-DPO}}$ recovers $f$-DPO as a special case
Consider $h$ corresponding to logistic regression, as discussed in Section A.5. We also consider the model ratio $R^{\text{$f$-DPO}}$, introduced in Section 3.4. Applying Eqs. 21 and 22 to $\mathcal{L}^{h}_{\textrm{BPO}}(R^{\text{$f$-DPO}}_{\theta};p_{\textrm{data}})$ yields $\mathcal{L_{\textrm{$f$-DPO}}}(\pi_{\theta};\pi_{\textrm{ref}},p_{\textrm{data}})$ in Eq. 23 as follows:
@@ B.1 $f$-DPO [54]
The related work $f$-DPO extends the RLHF formulation in Eq. 2 to $f$-divergence regularization:
@@
Optimality The optimal policy $\pi^{\textrm{$f$}}_{\theta^{*}}$ that minimizes $\mathcal{L}_{\textrm{$f$-RLHF}}$ or $\mathcal{L_{\textrm{$f$-DPO}}}$ is given by
@@
where $Z_{f}({\mathbf{x}})=\sum_{{\mathbf{y}}}\pi_{\text{ref}}({\mathbf{y}}|{\mathbf{x}})(f^{\prime})^{-1}\left(\frac{1}{\beta}r_{\phi^{*}}({\mathbf{x}},{\mathbf{y}})\right)$ is the partition function. Importantly, $\pi^{\textrm{$f$}}_{\theta^{*}}$ varies with the choice of $f$, indicating that this extension does not preserve the optimality of the original DPO. This is in contrast to BPO, which preserves the optimal solution as DPO for any general choice of $h$, as proved in Theorem 2.
@@ B.2 $f$-PO [59]
Formulation that preserves optimality The traditional generative modeling studies [38, 60] formulate to minimize $f$-divergence between model distribution $\pi_{\theta}$ and target distribution $\pi_{\theta}^{*}$ as $D_{f}(\pi_{\theta}||\pi_{\theta^{*}})$, as described in the third row of Table 1. The $f$-PO slightly re-defines the model and target distribution as:
@@
and formulate the matching problem as $D_{f}(\hat{\pi}_{\theta}||\hat{\pi}^{*})$, which still guarantees $\pi_{\theta}$ converges at $\pi_{\theta}^{*}$ at the optimum. We denote this property of the objective as O in Table 1. The $f$-divergence objective becomes:
@@
Simplified $f$-PO In the case of using pairwise preference datasets, an approximate version was also proposed to remove the additional computational burden mentioned above. If we perform Monte Carlo (MC) estimation using only two samples and set the positive reward function value to $1$ and the negative to 00, the objective becomes as follows (the original paper also presents a formulation that includes label smoothing, but we omit it for simplicity):
@@
Orthogonal extension of BPO to simplified $f$-PO. As discussed in Section 3.4, losses that can be expressed as logistic regression can be extended using BPO. Similarly, some instances of simplified $f$-PO can also be applied within the BPO framework by interpreting $f$-PO as logistic regression. The objective $\mathcal{L}_{\textrm{$f$-PO}}$ in Eq. 24 is equivalent to
@@
and $\mathcal{L}^{\text{h}}_{\text{BPO}}(R_{\theta}^{f\text{-PO}};p_{\text{data}})$ recovers Eq. 25 with $h$ corresponding to the logistic regression, while allowing further generalization of the loss by choosing different forms of $h$.
B  Formal discussions of related work (continued)
@@ B.3 SimPO [36]
$\mathcal{L}^{h}_{\textrm{BPO}}(R_{\theta}^{\text{SimPO}};p_{\textrm{data}})$ generalizes the $\mathcal{L_{\textrm{SimPO}}}$ with the choice of $h$ similar to Section 3.4. Unlike the model ratios $R_{\theta}$ in Eq. 8 or $R^{f{\text{-DPO}}}_{\theta}$ in Eq. 14, the value at initialization is not 1 but is close to exp$(\gamma)$. To match the gradient scale between the logistic regression loss and the SBA loss, we accordingly adjusted the scaling factor $s$ to match the gradient scale.
@@ B.4 DDRO [66]
where $p^{w}_{\text{data}}$ denotes the preferred data distribution and $p^{l}_{\text{data}}$ the dispreferred one. DDRO aims to match the policy model $\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})$ to the preferred distribution $p^{w}_{\text{data}}({\mathbf{y}}_{w}|{\mathbf{x}})$, reflecting a different notion of optimality from ours. In DDRO, although the dispreferred dataset is used in the loss computation, it does not affect the definition of the optimal policy. As shown in Proposition 1, our definition of the optimal policy reflects both the preferred and dispreferred distribution with point-wise human preference. The resulting objective of DDRO is defined as:
@@ C.1 Training algorithm
Algorithm 1 describes the detailed algorithm for implementing the general BPO objective with any valid function $h$. The main difference from the original DPO lies in Algorithm 1, where the loss is computed. Code LABEL:lst:bpo provides a PyTorch implementation of Algorithms 1 and 1.
@@ C.2 Hyperparameter setups
Dialogue generation task We follow the setup described in the official DPO [41] repository [^1] https://github.com/eric-mitchell/direct-preference-optimization for the dialogue generation task. We use Pythia-2.8B [5] [^2] https://huggingface.co/EleutherAI/pythia-2.8b as the pre-trained LLM, and perform supervised fine-tuning (SFT) for one epoch using the 161k prompt–preferred response pairs from the training set of the Anthropic helpful and harmless (HH) dataset [3] [^3] https://huggingface.co/datasets/Anthropic/hh-rlhf. SFT is conducted with a batch size of 64 using the RMSprop optimizer with a learning rate of 5e-7. We use the resulting checkpoint as a reference model for all experiments reported in Figure 1, Table 3, Table 4, Figure 4, and Figure 5 in the main paper.
@@
Preference optimization starts from the SFT model and is performed for one epoch using the training set of the HH dataset. All experiments reported in Figure 1, Table 3, Table 4, Figure 4, and Figure 5 share the same default training configuration: $\beta=0.1$, a batch size of 64, and the RMSprop optimizer with a learning rate of 5e-7. For $f$-PO [59], we apply label smoothing with $\epsilon=1\times 10^{-3}$, as recommended in their paper, to improve empirical stability. We also observe that BPO sometimes suffers from large gradient magnitudes when the ratio value becomes too small. To address this, we clip $R_{\theta}$ values to be no smaller than 0.01, which improves training stability. For generations, we use a sampling temperature of 1.0, which generally yields the best win rate on the HH dataset.
@@
Summarization task We follow the same training configuration as the dialogue task using a public SFT model [^4] https://huggingface.co/CarperAI/openai_summarize_tldr_sft, except that we set $\beta$ to 0.5 for all methods, as specified in the DPO paper. We perform preference optimization for one epoch on the 93k training subset of the comparison version of the TL;DR summarization dataset [^5] https://huggingface.co/datasets/openai/summarize_from_feedback. The same setting is used for measuring the performance reported in Table 5 and Figure 6. For BPO, we use SBA loss with $\lambda=-0.5$ and an $R_{\theta}$ clipping value of 0.025.
@@
External benchmarking task For Mistral-7B [23] [^6] https://huggingface.co/alignment-handbook/zephyr-7b-sft-full, we perform preference optimization for one epoch on 61k samples from the UltraFeedback [^7] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized dataset. For the DPO model ratio, we use a batch size of 64 and a learning rate of 8e-7. SBA loss is applied with $\lambda=-0.5$, and $R_{\theta}$ is clipped to be no smaller than 0.003. For the SimPO model ratio, we use the same setup as SimPO, applying SBA loss with $\lambda=-0.5$ and clipping $R_{\theta}$ to lie in the range [0.01, 30]. We interpolate between the SBA loss and the logistic regression loss, where the resulting objective still falls under the BPO framework. For Llama-3-8B-Instruct [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, we perform preference optimization for one epoch on the modified UltraFeedback [^9] https://huggingface.co/datasets/princeton-nlp/llama3-ultrafeedback-armorm dataset. We use the same SimPO model ratio with $\lambda=-0.5$ and clip $R_{\theta}$ to the range [0.01, 100]. For decoding parameters and templates, we follow the SimPO repository [^10] https://github.com/princeton-nlp/SimPO and ensure compatibility with the benchmark library configurations, including alpaca-eval==0.6.2 and vllm==0.5.4.
@@ C.4 Evaluation metrics
This section outlines the evaluation metrics used for the dialogue generation and summarization tasks in Section 4.1. Model performance is evaluated on a held-out test set, using 100 randomly sampled examples.
Predictive entropy This metric evaluates the diversity of model responses by directly reflecting the probabilities assigned to each generated token. For each prompt, we sample five responses and compute the predictive entropy using the implementation from $f$-DPO [54] [^11] https://github.com/alecwangcq/f-divergence-dpo/blob/main/metrics/imdb/imdb_eval_metrics.py.
Self-Bleu This metric measures the diversity of generated sentences. For each prompt, we generate five samples and compute the BLEU [64] score for each pair of responses, then take the average. The final score is obtained by averaging over all prompts.
Disctinct-1 This metric assesses the lexical diversity of a single response by computing the ratio of unique unigrams to the total number of unigrams [27]. For each prompt, we generate one sample and calculate this value. A higher score indicates greater lexical diversity within the response.
GPT-4 win rate We adopt the prompt template proposed in DPO [41], as shown below. To mitigate potential position bias, we evaluate each pair twice using GPT-4 as the judge model, swapping the positions of A and B.
@@ D.1 Additional $\lambda$ ablations
Figure 8 presents alternative metric results corresponding to Figure 5, and exhibits a performance sweet spot at a similar value of $\lambda$. Table 7 extends Table 4 by showing SBA results across different $\lambda$ values for each $f$. While Table 4 reports results with fixed $\lambda$ values ($\lambda=0.5$ for JS and $\lambda=0$ for FKL), Table 7 provides a broader view by sweeping over multiple values. Although the optimal $\lambda$ tends to differ slightly from the DPO model ratio case, SBA often outperforms the vanilla LR baseline.
@@
Figure 8: Additional ablation studies on the effect of $\lambda$ in SBA, using Anthropic-HH with Pythia-2.8B.

### v2

@@ A.1 Proof of Proposition˜1
From the Bradley-Terry model we discussed in Eq.˜5, the preference data distribution $p_{\text{data}}({\mathbf{y}}_{w}\succ{\mathbf{y}}_{l}|{\mathbf{x}})$ can be expressed as
@@ A.2 Proof of Theorem˜2
Let $\operatorname*{arg\,min}_{\pi_{\theta}}D_{h}(R_{\text{data}}||R_{\theta})=\pi_{\hat{\theta}}$. We have followings from Eqs.˜6 and 9:
@@
From the property of Bregman divergence [10], $B_{h}\left(R_{\text{data}}({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})||R_{\theta}({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})\right)$ achieves its minimum value of 0 if and only if $R_{\text{data}}({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})=R_{\theta}({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$ for a given point $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$. Under the assumption that the data distribution is fully supported, the minimizer $\pi_{\hat{\theta}}$ is guaranteed to satisfy $R_{\text{data}}=R_{\hat{\theta}}$ for all $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$. From the definition of $R_{\text{data}}$ and $R_{\theta}$ in Eq.˜8, we obtain the following:
@@
From the definition of $\pi^{*}$ in Proposition˜1, we obtain the following relation, which holds for all points $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$.
@@ A.5 BPO recovers DPO as a special case
This section shows that BPO includes DPO as a special case as discussed in Section˜3.3. Consider $h(R)=\frac{R\log{R}-(1+R)\log{(1+R)}}{2}$ which corresponds to the logistic regression in Table˜2, which has a derivative function as $h^{\prime}(R)=\frac{\log{R}-\log{(1+R)}}{2}$ and the expression inside the expectation operator of $\mathcal{L}^{h}_{\textrm{BPO}}(R_{\theta};p_{\textrm{data}})$ in Eq.˜10 becomes:
@@ A.6 BPO with $R_{\theta}^{f\text{-DPO}}$ recovers $f$-DPO as a special case
Consider $h$ corresponding to logistic regression, as discussed in Section˜A.5. We also consider the model ratio $R^{\text{$f$-DPO}}$, introduced in Section˜3.4. Applying Eqs.˜21 and 22 to $\mathcal{L}^{h}_{\textrm{BPO}}(R^{\text{$f$-DPO}}_{\theta};p_{\textrm{data}})$ yields $\mathcal{L_{\textrm{$f$-DPO}}}(\pi_{\theta};\pi_{\textrm{ref}},p_{\textrm{data}})$ in Eq.˜23 as follows:
@@ B.1 $f$-DPO [62]
The related work $f$-DPO extends the RLHF formulation in Eq.˜2 to $f$-divergence regularization:
@@
Optimality: The optimal policy $\pi^{\textrm{$f$}}_{\theta^{*}}$ that minimizes $\mathcal{L}_{\textrm{$f$-RLHF}}$ or $\mathcal{L_{\textrm{$f$-DPO}}}$ is given by
@@
where $Z_{f}({\mathbf{x}})=\sum_{{\mathbf{y}}}\pi_{\text{ref}}({\mathbf{y}}|{\mathbf{x}})(f^{\prime})^{-1}\left(\frac{1}{\beta}r_{\phi^{*}}({\mathbf{x}},{\mathbf{y}})\right)$ is the partition function. Importantly, $\pi^{\textrm{$f$}}_{\theta^{*}}$ varies with the choice of $f$, indicating that this extension does not preserve the optimality of the original DPO. This is in contrast to BPO, which preserves the optimal solution as DPO for any general choice of $h$, as proved in Theorem˜2.
@@ B.2 $f$-PO [67]
Formulation that preserves optimality: The traditional generative modeling studies [46, 68] formulate to minimize $f$-divergence between model distribution $\pi_{\theta}$ and target distribution $\pi_{\theta}^{*}$ as $D_{f}(\pi_{\theta}||\pi_{\theta^{*}})$, as described in the third row of Table˜1. The $f$-PO slightly re-defines the model and target distribution as:
@@
and formulate the matching problem as $D_{f}(\hat{\pi}_{\theta}||\hat{\pi}^{*})$, which still guarantees $\pi_{\theta}$ converges at $\pi_{\theta}^{*}$ at the optimum. We denote this property of the objective as O in Table˜1. The $f$-divergence objective becomes:
@@
Simplified $f$-PO: In the case of using pairwise preference datasets, an approximate version was also proposed to remove the additional computational burden mentioned above. If we perform Monte Carlo (MC) estimation using only two samples and set the positive reward function value to $1$ and the negative to 0, the objective becomes as follows (the original paper also presents a formulation that includes label smoothing, but we omit it for simplicity):
@@
Orthogonal extension of BPO to simplified $f$-PO: As discussed in Section˜3.4, losses that can be expressed as logistic regression can be extended using BPO. Similarly, some instances of simplified $f$-PO can also be applied within the BPO framework by interpreting $f$-PO as logistic regression. The objective $\mathcal{L}_{\textrm{$f$-PO}}$ in Eq.˜24 is equivalent to
@@
and $\mathcal{L}^{\text{h}}_{\text{BPO}}(R_{\theta}^{f\text{-PO}};p_{\text{data}})$ recovers Eq.˜25 with $h$ corresponding to the logistic regression, while allowing further generalization of the loss by choosing different forms of $h$.
@@ B.3 SimPO [43]
[…]
@@ B.4 DDRO [22]
where $p^{w}_{\text{data}}$ denotes the preferred data distribution and $p^{l}_{\text{data}}$ the dispreferred one. DDRO aims to match the policy model $\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})$ to the preferred distribution $p^{w}_{\text{data}}({\mathbf{y}}_{w}|{\mathbf{x}})$, reflecting a different notion of optimality from ours. In DDRO, although the dispreferred dataset is used in the loss computation, it does not affect the definition of the optimal policy. As shown in Proposition˜1, our definition of the optimal policy reflects both the preferred and dispreferred distribution with point-wise human preference. The resulting objective of DDRO is defined as:
@@ C.1 Training algorithm
Algorithm˜1 describes the detailed algorithm for implementing the general BPO objective with any valid function $h$. The main difference from the original DPO lies in Algorithm˜1, where the loss is computed. Code LABEL:lst:bpo provides a PyTorch implementation of Algorithms˜1 and 1.
@@ C.2 Hyperparameter setups
Dialogue generation task: We follow the setup described in the official DPO [49] repository [^1] https://github.com/eric-mitchell/direct-preference-optimization for the dialogue generation task. We use Pythia-2.8B [8] [^2] https://huggingface.co/EleutherAI/pythia-2.8b as the pre-trained LLM, and perform supervised fine-tuning (SFT) for one epoch using the 161k prompt–preferred response pairs from the training set of the Anthropic helpful and harmless (HH) dataset [6] [^3] https://huggingface.co/datasets/Anthropic/hh-rlhf. SFT is conducted with a batch size of 64 using the RMSprop optimizer with a learning rate of 5e-7. We use the resulting checkpoint as a reference model for all experiments reported in Figure˜1, Table˜3, Table˜4, Figure˜4, and Figure˜5 in the main paper.
@@
Preference optimization starts from the SFT model and is performed for one epoch using the training set of the HH dataset. All experiments reported in Figure˜1, Table˜3, Table˜4, Figure˜4, and Figure˜5 share the same default training configuration: $\beta=0.1$, a batch size of 64, and the RMSprop optimizer with a learning rate of 5e-7. For $f$-PO [67], we apply label smoothing with $\epsilon=1\times 10^{-3}$, as recommended in their paper, to improve empirical stability. We also observe that BPO sometimes suffers from large gradient magnitudes when the ratio value becomes too small. To address this, we clip $R_{\theta}$ values to be no smaller than 0.01, which improves training stability. For generations, we use a sampling temperature of 1.0, which generally yields the best win rate on the HH dataset.
@@
Summarization task: We follow the same training configuration as the dialogue task using a public SFT model [^4] https://huggingface.co/CarperAI/openai_summarize_tldr_sft, except that we set $\beta$ to 0.5 for all methods, as specified in the DPO paper. We perform preference optimization for one epoch on the 93k training subset of the comparison version of the TL;DR summarization dataset [^5] https://huggingface.co/datasets/openai/summarize_from_feedback. The same setting is used for measuring the performance reported in Table˜5 and Figure˜6. For BPO, we use SBA loss with $\lambda=-0.5$ and an $R_{\theta}$ clipping value of 0.025.
@@
External benchmarking task: For Mistral-7B-Base [28] [^6] https://huggingface.co/alignment-handbook/zephyr-7b-sft-full, we perform preference optimization for one epoch on 61k samples from the UltraFeedback [^7] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized dataset. For the DPO model ratio, we use a batch size of 64 and a learning rate of 8e-7. SBA loss is applied with $\lambda=-0.5$, and $R_{\theta}$ is clipped to be no smaller than 0.003. For the SimPO model ratio, we use the same setup as SimPO, applying SBA loss with $\lambda=-0.5$ and clipping $R_{\theta}$ to lie in the range [0.01, 30]. We interpolate between the SBA loss (0.3) and the logistic regression loss (0.7), where the resulting objective still falls under the BPO framework. For Llama-3-8B-Instruct [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, we perform preference optimization for one epoch on the modified UltraFeedback [^9] https://huggingface.co/datasets/princeton-nlp/llama3-ultrafeedback-armorm dataset. We use the same SimPO model ratio with $\lambda=-0.5$ and clip $R_{\theta}$ to the range [0.01, 100]. We interpolate between the SBA loss (0.05) and the logistic regression loss (0.95). For Llama-3-8B-Base, we use the same SimPO model ratio with $\lambda=-0.5$ and clip $R_{\theta}$ to the range [0.01, 30]. We interpolate between the SBA loss (0.5) and the logistic regression loss (0.5). For decoding parameters and templates, we follow the SimPO repository [^10] https://github.com/princeton-nlp/SimPO and ensure compatibility with the benchmark library configurations, including alpaca-eval==0.6.2 and vllm==0.5.4.
@@ C.4 Evaluation metrics
This section outlines the evaluation metrics used for the dialogue generation and summarization tasks in Section˜4.1. Model performance is evaluated on a held-out test set, using 100 randomly sampled examples.
Predictive entropy: This metric evaluates the diversity of model responses by directly reflecting the probabilities assigned to each generated token. For each prompt, we sample five responses and compute the predictive entropy using the implementation from $f$-DPO [62] [^11] https://github.com/alecwangcq/f-divergence-dpo/blob/main/metrics/imdb/imdb_eval_metrics.py.
Self-Bleu: This metric measures the diversity of generated sentences. For each prompt, we generate five samples and compute the BLEU [73] score for each pair of responses, then take the average. The final score is obtained by averaging over all prompts.
Disctinct-1: This metric assesses the lexical diversity of a single response by computing the ratio of unique unigrams to the total number of unigrams [34]. For each prompt, we generate one sample and calculate this value. A higher score indicates greater lexical diversity within the response.
GPT-4 win rate: We adopt the prompt template proposed in DPO [49], as shown below. To mitigate potential position bias, we evaluate each pair twice using GPT-4 as the judge model, swapping the positions of A and B.
@@ D.1 Additional $\lambda$ ablations
Figure˜8 presents alternative metric results corresponding to Figure˜5, and exhibits a performance sweet spot at a similar value of $\lambda$. Table˜7 extends Table˜4 by showing SBA results across different $\lambda$ values for each $f$. While Table˜4 reports results with fixed $\lambda$ values ($\lambda=0.5$ for JS and $\lambda=0$ for FKL), Table˜7 provides a broader view by sweeping over multiple values. Although the optimal $\lambda$ tends to differ slightly from the DPO model ratio case, SBA often outperforms the vanilla LR baseline.
@@
Figure 8: Additional ablation studies on the effect of $\lambda$ in SBA, using Anthropic-$H$ with Pythia-2.8B.